### v1

# BuildingView: Constructing Urban Building Exteriors Databases with Street View Imagery and Multimodal Large Language Model

Abstract.
Urban Building Exteriors are increasingly important in urban analytics, driven by advancements in Street View Imagery and its integration with urban research. Multimodal Large Language Models (LLMs) offer powerful tools for urban annotation, enabling deeper insights into urban environments. However, challenges remain in creating accurate and detailed urban building exterior databases, identifying critical indicators for energy efficiency, environmental sustainability, and human-centric design, and systematically organizing these indicators. To address these challenges, we propose BuildingView, a novel approach that integrates high-resolution visual data from Google Street View with spatial information from OpenStreetMap via the Overpass API. This research improves the accuracy of urban building exterior data, identifies key sustainability and design indicators, and develops a framework for their extraction and categorization. Our methodology includes a systematic literature review, building and Street View sampling, and annotation using the ChatGPT-4O API. The resulting database, validated with data from New York City, Amsterdam, and Singapore, provides a comprehensive tool for urban studies, supporting informed decision-making in urban planning, architectural design, and environmental policy. The code for BuildingView is available at https://github.com/Jasper0122/BuildingView.
Building exteriors, Street-View imagery, Multimodal large language model, GeoAI, Urban analytics
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Urban environments are increasingly becoming the focus of interdisciplinary research, where the integration of spatial and visual data is essential for understanding various aspects of urban life. Within this context, the study of building exteriors has gained significant attention due to its direct impact on critical areas such as energy efficiency(parking-lots, ), environmental sustainability(low-carbon, ), and human-centric urban design(human-perception, ). Building exteriors serve not only as the physical interface between a structure and its environment but also play a crucial role in determining the overall energy performance and aesthetic appeal of urban spaces. Street view imagery, in particular, has emerged as a powerful tool in this domain, offering a unique and detailed perspective on building exteriors that traditional remote sensing methods often overlook(liang2024evaluating, ). Unlike remote sensing imagery, which provides a top-down view of urban landscapes, street view imagery captures buildings at eye level, revealing intricate details of their spatial attributes, materials, and architectural styles(biljecki2021street, ). This granular level of detail is vital for assessing the impact of building exteriors on factors such as urban heat islands, energy consumption, and the overall aesthetic coherence of neighborhoods. While street view imagery offers these advantages, its potential remains underutilized in urban research, particularly when it comes to integrating this visual data with other spatial datasets to provide a comprehensive understanding of building exteriors and their broader urban implications. This highlights the growing need to harness street view imagery more effectively to address the multifaceted challenges of modern urban environments.
Through a systematic search of the urban building exteriors research, it is evident that there is a significant gap in the availability of comprehensive databases that integrate street view imagery with detailed building exterior information. Existing studies tend to focus on either the spatial characteristics of buildings—such as their geographic positioning and interactions with surrounding infrastructure—or the aesthetic and material aspects, including innovations in architectural design and material usage(building-style, )(material, ). However, these perspectives are rarely combined to offer a full understanding of urban building exteriors. The lack of integrated databases that unify spatial data with visual and material information limits the ability to conduct in-depth analyses necessary for informed decision-making in urban planning, architectural design, and environmental policy. Additionally, while recent advancements have leveraged street view imagery for various urban analyses, they primarily focus on environmental factors or general urban features rather than systematically extracting and organizing building exterior information(rundle2011using, ). The integration of large language models with street view imagery for detailed urban analysis remains in its nascent stages, highlighting the need for a more integrated approach to understanding urban building exteriors.
To address these knowledge deficits, we propose BuildingView, an innovative approach that integrates Street View imagery with multimodal large language models to create a comprehensive urban building exteriors database. By leveraging high-resolution visual data from Google Street View, combined with spatial and structural data obtained through the Overpass API and OpenStreetMap, this research sets a new standard for the integration of visual and spatial data in urban informatics, significantly advancing the field. Our key contributions include: (1) the systematic collection and organization of urban building exteriors indicators, providing a comprehensive understanding of critical factors such as energy efficiency, environmental sustainability, and human-centric design. (2) the development of BuildingView, a reusable tool that utilizes Street View imagery and multimodal large language models to construct detailed urban building exterior databases for any urban area globally. (3) the construction of extensive urban building exterior databases for New York, Amsterdam and Singapore, enabling more informed decision-making in urban planning, architectural design, and environmental policy.
2. Related Work
2.1. Indicators of Building Exteriors Indicators
In the realm of urban informatics and geoscience, the development of spatiotemporal databases pertaining to buildings has garnered significant attention. Researchers within the disciplines of urban planning and geography typically concentrate on the spatial characteristics of buildings, such as their geographic positioning and the dynamics of their interaction with adjacent infrastructure. For instance, the architectural and urban planning processes must meticulously consider the spatial interplay between buildings and their adjacent parking facilities to optimize energy efficiency (parking-lots, ).In contrast, architects and designers are more inclined to examine the intrinsic elements of buildings, including materials and aesthetic design. Recently, the architectural style and the materials used for the exterior of buildings are continually subject to innovation and enhancement (building-style, ; material, ).
The academic discourse has increasingly pivoted towards the examination of the connections between buildings and other indicators, with a particular emphasis on their impact on energy consumption and environmental sustainability. Especially, researchers are actively exploring ways to harness more energy from natural sources through building design, thereby reducing the energy costs associated with buildings. Additionally, they are working to mitigate the thermal impacts caused by the urban heat island effect, which is exacerbated by the heat generated within urban environments (energy, ). Concurrently, there is a burgeoning interest in the human-centric aspects of building design, particularly how it fosters engagement and interaction with the public from a participatory perspective (human-perception, ).
While numerous open-source databases exist that cater to specific domains of building information, they are often limited to a single aspect of building data. The foundational building attribute knowledge within these databases is typically derived from field surveys or footprint data extracted from remote sensing imagery (chen2022large, ). To date, no open-source database has been developed that integrates street view imagery to provide a comprehensive, visually-driven repository encompassing the spatial. Thus, we want to build a spatial database with fundamental, and external surface properties of buildings, with a focus on their relevance to energy, environmental, and human-centric dimensions.
2.2. SVI and LLMs in Urban Building Analysis
The advent of street view imagery (SVI) has revolutionized the way urban environments are analyzed, offering a rich and dynamic geospatial data source that rivals traditional remote sensing methods such as satellite imagery (biljecki2021street, ). The comprehensive visual data captured by SVI allows for detailed analysis and understanding of urban environments at the street level, providing critical insights into various urban characteristics and phenomena.
Previous research has demonstrated the utility of SVI in urban studies by integrating it with other urban data sources such as street networks, building information, demographic and socioeconomic data, and survey responses. For instance, SVI has been utilized to audit neighborhood environments (rundle2011using, ), and analyze the effects of green view indices on pedestrian activity(ki2021analyzing, ). Additionally, studies have explored the use of SVI and machine learning to measure the built environment’s impact on crime (hipp2021measuring, ), as well as investigating the relationship between neighborhood walkability factors and walking behaviors using big data approaches (koo2022neighborhood, ). Furthermore, SVI has been applied at a global scale, as seen in the creation of extensive datasets that cover millions of street-level images across hundreds of cities, facilitating large-scale urban science and analytics(hou2024global, ).
However, to date, there has been no significant work that focuses on using SVI in conjunction with LLMs to construct a comprehensive database of urban building exteriors. The proposed BuildingView project aims to fill this gap by harnessing the combined power of SVI and multimodal LLMs to create an extensive and detailed urban building exteriors database, which can serve as a valuable resource for a wide range of urban studies, including architectural analysis, urban planning, and environmental assessment.
3. Method
In this study, we introduce BuildingView, a novel framework for constructing urban building exteriors databases using Street View imagery and multimodal large language models. The research framework is divided into three primary steps, as illustrated in Figure 1: (1) a systematic literature review of building exterior indicators; (2) sampling of buildings and street view imagery using the Overpass and Google Street View APIs to gather data points; and (3) annotation of street view images using the ChatGPT-4O API to create the Urban Building Exteriors Database. Further details on each of these steps are provided below.
Figure 1. The Workflow of BuildingView
[…]
A comprehensive workflow for analyzing building exteriors using a systematic literature review, sampling of buildings and street view imagery, and annotation with a multimodal large language model, integrating multiple APIs and exporting data in various spatial formats.
3.1. Systematic Literature Review of Building Exteriors Indicators
Global building morphology indicators database has concluded many indicators for urban data science (global-indicators, ). However, a database of buildings directly covering energy, green, and human-centred buildings has not yet been created. Our goal is to establish a building database that encompasses the majority of indicators of interest to researchers. This database will include comprehensive spatial information on buildings and knowledge extracted from street view images through advanced vision analytics and vision-based reasoning techniques.
For that, we have done a systematic literature review on building exteriors Figure 2. The initial step is to identify emerging topics concerning building exteriors over the past five years. To achieve this, we conduct a keyword search on ”building envelope,” ”exterior of building,” and ”building exterior” within the Web of Science database. Our search yields 6,28562856,2856 , 285 relevant scholarly works. There is a growing discussion about the intelligent design of external surfaces of buildings. The integration of innovative materials and intelligent technologies has the potential to significantly enhance the comfort of indoor environments (wang2024advancements, ).
Figure 2. Literature Review of Building Exteriors
[…]
Literature Review of Building Exteriors on Web of Science.
In addition to fundamental building information and its spatial attributes, we have categorized the concept of ”smart building exterior design” into three primary areas that are most frequently discussed by scholars: ”human,” ”energy,” and ”green.” Incorporating these keywords into our search criteria, we have identified a total of 4,68346834,6834 , 683 relevant works, with the distribution being $544544$ for ”human,” 4,11541154,1154 , 115 for ”energy,” and $523523$ for ”green.”
Given our focus on indicators that can be discerned through visual analysis, we have narrowed down the total literature on building exteriors to $5454$ relevant pieces. Within this subset, only $3939$ pieces specifically address the hot topic of human-centric, energy fuel and engineering environmental fields. We reviewed these works and we found that among these, the topic of glazing has garnered significant attention. Specific indicators such as glazing type, colour, and Window-Wall Ratio (WWR) are recognized as crucial factors influencing energy absorption (WWR, ; WWR-and-type, ). Outdoor air conditioners, as significant contributors to heat, and their placement methods are also central topics in discussions about energy conservation (aircon-position, ). The rooftop, as the uppermost exterior of the building, plays a pivotal role in capturing solar energy. Its design and materials can significantly influence the building’s energy efficiency by absorbing and utilizing solar radiation (energy-saving, ). From a human-centric perspective, building styles exhibit considerable variation, influenced by factors such as culture, climate, and history (ng2020perception, ). They are commonly recognized as falling into categories such as Historical, Somewhat Historical, No Significant Style, Somewhat Modern, and Modern (human-perception, ). Vertical greenery not only contributes to aesthetics but also has a positive impact on the thermal insulation of building facades (campiotti2022vertical, ). Meanwhile, the surrounding infrastructure, such as transport stations and parking lots, can significantly influence the design of building exteriors and is, in turn, affected by them (transport, ; parking-lots, ). Consequently, we will also consider these elements.
Subsequently, we have identified and summarized the key indicators that researchers find intriguing, which are presented in Table. 1. The indicators, which can be derived through vision-based reasoning from street view maps, are predominantly self-explanatory. However, some may not be readily distinguishable by AI technology, and thus, we provide additional explanations for a few that require further clarification. We have considered a range of indicators for our research on building exteriors, encompassing basic building information such as architectural style, building type, and relative location, as well as characteristics like colour. We have also included energy-saving indicators, environmental friendliness indicators—such as Window-Wall Ratio (WWR) and building exterior materials, vertical vegetation type, and human-centric indicators visible on the exteriors, including streetscape perception scores and exterior complexity, etc. Additionally, we have taken into account indicators of the surrounding environment that may influence the design of building exteriors. In parallel, we have integrated useful information for 3D building analysis, including details about neighbouring buildings and greening conditions. Subsequently, we have developed systematic and comprehensive prompts to assist ChatGPT in extracting these indicators from Street View images, aiming to obtain an accurate, detailed, and structured description of the attributes of our building exteriors to fulfill our global spatial database about building exteriors.
Table 1. Building Exteriors Indicators
Indicator Value Example Explanation
Architectural style Haussman, Neoclassical, Renaissance, Modernism, Others.
Building type single-family houses, multiple-family houses, non-residential buildings based on the number of units
Relative location on the surface, in the air, underground, across the surface relative location between building property and the ground surface
Colour red, grey, light brown, unknown. the colour of the building exteriors
Floor-to Floor height 3m, approximately 2m, unknown. inferences based on a prior knowledge
Number of Doors and Windows 3 doors and approximately 2 windows, unknown.
Number of Floors 2, approximately 2, unknown. inferences based on a prior knowledge
WWR Approximately 0.25, 30%, unknown. the ratio of the total area of windows to the total wall area
Glazing type single, double, or triple, unknown types refer to the number of glass layers in windows
Window colour grey, light brown, unknown. the colour of the exterior glazing of windows
Material stone, glass product, metallic materials, or indistinguishable due to exterior paint the material of building exteriors
Classification of the building materials alternative materials, natural materials, or secondary raw materials the classification based on building exterior materials
Number of vertices
Vertical greenery type panel type, mini planter, cage system box, unknown the type of vertical greenery system on the building exteriors
Roof type lightweight, green, photovoltaic, vents, rubber the classification of roof types from an energy-saving perspective
Number of air-con units 3, approximately 3, unknown number of outdoor air conditioning units
air-con placement type horizontal, vertical placement type of outdoor air-con units
Street type residential street, local street, etc.
Neighbouring buildings 3 similar buildings, none, etc.
Greening conditions 3 trees, 3 grasslands, unknown, etc.
Street facilities 3 roads, 3 parking lots, unknown, etc.
Public transport 3 subway stations, 3 bus stops, none, unknown, etc.
Human perceptual ratings complex, original, ordered, pleasing, boring
building style Historical, somewhat historical, no significant style, somewhat modern, modern Style of the building from a human-centric perspective
Exterior complexity complex, moderate, simple inferences based on a prior knowledge
Streetscape perception scores Safer, wealthier, livelier, more beautiful, more depressing, more boring
3.2. Sampling of Buildings and Street View Imagery
We begin by using the Nominatim API to fetch city names and their corresponding countries based on a query provided through the command line. Nominatim, an open-source geocoding service, outputs a list of cities and countries that match the query, allowing for accurate identification of the desired city and country(clemens2015geocoding, ). This makes it easier to directly locate the corresponding bounding box by simply inputting the city and country names, streamlining the data collection process.
For the building sampling, we employ the Overpass API to retrieve and save building data for the specified city and country. The Overpass API(olbricht2011overpass, ), which is used to query the OpenStreetMap (OSM) database, performs several tasks: it retrieves the bounding box coordinates for the given city and country, queries the OSM database to obtain building IDs and their coordinates within the bounding box, gathers additional details such as the address and height of each building, and saves the building data into a JSONL file, organized by building types.
Finally, to match the buildings with their corresponding street view exteriors, we use the Google Street View API to download images for the locations specified in the JSONL file. This process involves creating a directory to store the images based on the name of the JSONL file, reading the latitude and longitude coordinates from the file, and downloading images using the Google Street View API for each specified location. The images are then saved in the created directory. To ensure accuracy, the building ID coordinates are centered within a 30-meter buffer, and the closest street view image is selected for each building, saved in a $600\times 300600 × 300$ format.
3.3. Annotation with Multimodal Large Language Model
We annotate the street view imagery using OpenAI’s ChatGPT-4o, a state-of-the-art language model designed to generate human-like text based on tailored prompts(openai2023chatgpt, ). To ensure stability and efficiency in API requests, we implement a robust, multi-threaded framework capable of managing multiple API keys to optimize the distribution of API calls. The prompts for annotation are meticulously developed based on urban building exterior indicators identified during our preliminary research. To maintain flexibility and adaptability in the annotation process, we incorporate customizable configurations, allowing modifications to suit specific research objectives.
We further streamline the annotation process by automating key tasks, including image downloading, error handling, and data management. This automation framework facilitates the execution of tasks according to predefined parameters, manages failed download attempts, consolidates data into a unified dataset while eliminating duplicate records, and tracks processed data to prevent redundancy.
The indicators resulting from our annotation process are designed to be exportable in various spatial data formats, including CSV, Shapefile, and GeoJSON. Our accompanying documentation provides several examples for querying and exporting data into both tabular and geospatial formats, ensuring compatibility with a wide range of analytical tools.
To promote accessibility and facilitate the distribution of ready-to-use datasets, we compile and publish datasets for multiple cities as open data within a dedicated repository. This repository includes datasets for cities and countries with detailed mapping available in OpenStreetMap (OSM), and it continues to expand as we add new locations. The ongoing development of this repository reflects our commitment to providing comprehensive and up-to-date urban building exterior data.
4. Experiment
4.1. Data
We construct an Urban Building Exteriors Database using Street View Imagery, theoretically applicable to any urban area with street views. For this project, we select three representative cities: New York City (NYC), Amsterdam, and Singapore. The distribution of data points for the three cities is shown in Figure3.
Figure 3. The Distribution of Sampling Points
[…]
The Distribution of Sampling Points.
Each city is chosen based on its unique architectural characteristics and urban landscape. New York City is renowned for its iconic skyscrapers and diverse architectural styles, ranging from historic brownstones to modern glass towers(lehan2023city, ). Amsterdam offers a distinct contrast with its picturesque canal houses, narrow buildings, and intricate facades, reflecting a historic and compact urban design(havinga2020heritage, ). Singapore, known for its modernist high-rise buildings and innovative architectural solutions, represents a dynamic and evolving urban environment that blends tradition with cutting-edge design(yan2022data, ).
Figure 4. Representative Street View Images from Selected Cities
[…]
This figure showcases representative street view images of building exteriors from New York, Amsterdam, and Singapore, highlighting the architectural diversity and urban environments captured in the dataset.
To collect building information, we employ the Overpass API, a powerful tool for querying OpenStreetMap(bennett2010openstreetmap, ) data, which enables us to compile comprehensive datasets for each city. We gather information on all fetched buildings within the three cities, totaling 1,382,71613827161,382,7161 , 382 , 716 buildings in NYC, 247,456247456247,456247 , 456 buildings in Amsterdam, and 298,787298787298,787298 , 787 buildings in Singapore. From these, we sample buildings and collect their building IDs, latitude and longitude coordinates, and address information. For the SVI, we utilize Google Street View Imagery(anguelov2010google, ) to capture panoramic images of individual building exteriors. To standardize the collection process, we center the building ID coordinates within a 30-meter buffer and select the closest street view image to each building. This method ensures consistency and precision in capturing building exteriors. The dataset comprises 8,13081308,1308 , 130 building images for NYC, 6,42264226,4226 , 422 building images for Singapore, and 7,75877587,7587 , 758 building images for Amsterdam. Additionally, sample street view images for each city have been included to provide visual context for the building exterior data collected, as shown in Figure 4.
4.2. Result and Analysis
We gather building exterior datasets from Amsterdam (7,75877587,7587 , 758), New York City (8,13081308,1308 , 130), and Singapore (6,42264226,4226 , 422), totaling 22,3102231022,31022 , 310 records. For most metrics, the effective generation rate exceeds 99.90%. However, the effective generation rate for streetscape perception scores is comparatively lower at 86.05%, primarily due to the inherent randomness in ChatGPT’s generation process and the complexities involved in regular expression matching.
To evaluate the model’s predictive performance, we select four key variables: Floor-to-Floor Height Numeric Only, Window-to-Wall Ratio (WWR), Vertices, and Tree Coverage. Using manual inspection as the benchmark, we randomly sample 200 instances from the three cities and calculate three key statistical indicators: Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and R². The results show that while predictive accuracy varies among different variables, the overall performance remains strong, as shown in Table 2. Notably, the R² values for Floor-to-Floor Height Numeric Only and WWR are $0.830360.83036$ and $0.746790.74679$, respectively. Although the Vertices variable presents certain challenges, its R² value of $0.697670.69767$ still indicates a reasonable level of predictive capability.
Table 2. Performance Metrics for Different Variables
Variable MAE RMSE R2
Floor-to-Floor Height 0.01724 0.09285 0.83036
WWR 0.01355 0.04130 0.74679
Vertices 0.82209 2.24292 0.69767
Tree 0.13462 0.50000 0.66215
Total 0.24687 0.71927 0.73424
In the accompanying figures, we select the following indicators from three perspectives—environment, energy, and human—to visualize the results: ’A’ represents the parking lot scale, ’B’ represents the tree coverage scale, and ’C’ represents the Window-to-Wall Ratio (WWR) scale, while the numbers 1, 2, and 3 correspond to Amsterdam, Singapore, and New York City, respectively. In Figure5, the color of the dots transitions from red to green, indicating the scale of the metrics from small to large. Deeper red signifies a smaller scale, while deeper green indicates a larger scale. The figure reveals that the parking lot scale has been most densely represented in red, particularly in the city centers; the tree coverage scale has been predominantly green, with darker hues; and the WWR scale has had the most intense green distribution.
Figure 5. The Visualization of Result: The Number of Parking Lots (A), Trees (b) and Window-to-Wall Ratio (c)
[…]
The Number of Parking Lots (A), Trees(b) and Window-to-Wall Ratio(c)
From the following word cloud, it is evident that the predominant architectural styles vary significantly across different cities. For example, in Figures A and C, ”somewhat modern” and ”modern somewhat” have been the most prevalent, whereas in Figure B, ”somewhat historical” dominates, with a relatively smaller presence in Figures A and C (see Figure6).
Figure 6. The Visualization of Result: Human Style Wordclouds
[…]
Human Style Wordclouds
5. Discussion
The building’s exterior has garnered considerable attention among researchers in the GIS domain and those engaged in urban and architectural studies. Concurrently, an increasing array of interdisciplinary professionals is examining the multifaceted influence that buildings exert on human life, the natural environment, and energy consumption. Conventional spatial databases often rely on survey data, with building attribute information typically confined to specific domains. Our conviction is that the academic community requires a comprehensive, global spatial database that documents the appearance of buildings, encompassing a wide array of significant locations.
Based on our literature review of building exteriors, we have categorized building exterior indicators across human-centric, energy, environmental, and basic information of buildings and initiated a spatial database that covers these indicators using the Large Language Model and computer vision. The characterization of building exteriors in our spatial database is grounded in the interpretations provided by ChatGPT-4o. While ChatGPT-4o’s proficiency in image comprehension is widely recognized, the absence of innate knowledge necessitates a methodical approach. To this end, we randomly select $200200$ street-level images from our database and undertake a meticulous manual review to ensure the precise depiction of building exteriors. Given that the human-centric assessments of building exteriors are inherently subjective, we have forgone verifying human-centric indicators.
We have identified and scrutinized four key indicators to evaluate the fidelity of our data: floor-to-floor height, window-to-wall ratio (WWR), the architectural vertices of the buildings, and the number of surrounding trees. Our manual review Table 2 reveals that the vision-based experiment has yielded robust results. Despite the challenges posed by occlusions inherent in street views, ChatGPT-4o has shown remarkable accuracy in correlating building heights, extrapolating door and window features, and assessing the immediate built environment. However, the number of building vertices—a three-dimensional characteristic—proved to be less accurate when relying solely on a single street view perspective. This limitation is attributed to the depth of field constraints, which can obscure the full dimensionality of the structures in question.
Despite the availability of numerous methods to construct a comprehensive global architectural spatial database, the vision-based approach significantly reduces the expenditure of human and material resources. Our experiments, while promising, reveal areas for enhancement:
  1. (1)
    The absence of street views in certain locales indicates that the creation of a complete global database of building appearances should integrate additional data sources, such as satellite imagery and expert survey data. Future endeavours could focus on developing a multi-source spatial database for building exteriors.
  2. (2)
    The single-view street view imagery presents challenges with occlusion and the limitations of a single perspective, affecting the accuracy of chatGPT-4o’s estimation of 3D building exterior attributes. Future improvements can involve leveraging images from various perspectives to generate more precise exterior descriptions.
  3. (3)
    Our current classification of building exteriors is subjective based on research trends. Moving forward, we aim to employ a more systematic text processing methodology, including optimizing chatGPT prompts, utilizing advanced NLP for text categorization, and incorporating additional databases to refine the accuracy of building exterior descriptions.
6. Conclusion
In this study, we have compiled a list of hot topics related to building exteriors for the last five years. Twenty-six metrics were summarised and categorised into disciplines. Our study presents a novel spatial database for building exteriors, demonstrating the feasibility of integrating AI and street view imagery to enhance the completeness of the database. The validation of our approach highlights its potential for efficiency and accuracy, setting a precedent for similar endeavours. Such a workflow can be replicated for global regions, resulting in a worldwide database of building exteriors.
Upon enhancing the building classification approach through the integration of multi-source datasets, we anticipate the development of a comprehensive and accurate global building exteriors database. This envisioned database will serve as a solid foundation, providing invaluable insights and data to researchers and industry experts alike.